OpenStudy (anonymous):

Solve the equation by factoring: (6x + 5)^2 = 10 Please help

OpenStudy (whpalmer4):

Expand it out, then move the 10 to the left side by subtracting 10 from both sides. Now can you factor it?

OpenStudy (stamp):

Evaluate (6x+5)(6x+5). Then set the expression equal to zero, and factor your expression to find the x values that satisfy the equation.

OpenStudy (anonymous):

square root both side first so; 6x+5=3.16, then find x value by moving 5 to the other side and divide 6 for both side.

OpenStudy (whpalmer4):

@stamp you'll get the wrong answer if you do so! the solutions to (6x+5)^2 = 0 are not the same as the solutions to (6x+5)^2 = 10!

OpenStudy (whpalmer4):

@eunsongpark That's not solving by factoring...

OpenStudy (stamp):

@whpalmer4 I am saying expand your square term, subtract 10 (thereby setting it equal to 0), simplify your new quadratic expression, then factor to obtain solution.

OpenStudy (whpalmer4):

@stamp okay, I misread your description. I agree that what you just said will work.

OpenStudy (anonymous):

@whpalmer4 oops! you're right. I didn't even read the question carefully;;

OpenStudy (anonymous):

thank you for your hardwork

OpenStudy (whpalmer4):

Of course, having insisted that we solve this by factoring, it's an irreducible polynomial!

OpenStudy (whpalmer4):

You could solve it with the quadratic formula, or by completing the square, or just taking the square root as @eunsongpark suggested (I'd leave it in symbolic form rather than writing an approximation for the square root of 10, however).